# Undergraduate Course Allocation

### v1

@@ 4 Discussion
This paper explores the many-to-many matching problem that arises in undergraduate course allocation. We allow courses to prioritize students based on factors such as year of study and department and design a deterministic allocation mechanism, the Pseudo-Market with Priorities (PMP) mechanism, that respects course priorities. The PMP mechanism extends the approximate competitive equilibrium from equal incomes mechanism of Budish (2011) to settings with course priorities and has the same, small worst-case bound on the market-clearing error. It also satisfies desirable approximate notions of stability, efficiency, fairness, and strategy-proofness. Our empirical findings demonstrate that the PMP mechanism outperforms the Random Serial Dictatorship with optimal course reserves, a mechanism resembling the one widely used in practice, in terms of student satisfaction and allocation fairness.
The benefits of introducing the PMP mechanism depend on university-specific factors. Universities with only a few popular courses will likely see small gains from implementing any fairer mechanism. In contrast, when many courses are overdemanded, the mechanisms used in practice often lead to significant envy among students at the same level of priority. The PMP mechanism will be most effective in these environments, as it guarantees fairness without compromising priority constraints or student satisfaction.
A university’s course priority structure also plays a considerable role in its potential gains from implementing the PMP mechanism. With priorities based on seniority and student department, the random serial dictatorship and deferred acceptance mechanisms heavily rely on tie-breaking to allocate courses, resulting in a large percentage of students with envy toward others in the same year of study and department. Our Theorem 4 guarantees that, with a small enough budget inequality, the PMP mechanism bounds envy among these students by a single course. In contrast, if schools employ finer priority structures, using factors such as units completed or grade point average, the PMP and RSD mechanisms are likely to result in similar course allocation outcomes.
@@
We conclude by discussing several aspects pertinent to the practical implementation of the PMP mechanism, drawing on previous experience with implementing the A-CEEI mechanism at Wharton and other business schools. The model in Section 1 assumes that students can report their preferences over all possible course schedules. In practice, with thousands of available courses, reporting preferences requires a language simple enough for students to use and expressive enough to elicit students’ actual preferences. The practical implementation of the A-CEEI mechanism in Budish et al. (2017) allows students to report cardinal values for individual courses and adjustments for pairs of courses. However, Budish and Kessler (2022) find evidence of preference reporting errors that negatively impacted student welfare at Wharton. Moreover, some adjustments to the preference reporting language of Budish et al. (2017) may be necessary to make this framework accessible to undergraduate students. For example, as credit limits play a large role in determining course schedules, students need to be able to indicate a maximum number of courses or credits to enroll in while reporting their preferences. Importantly, the deployment of recent machine learning-based techniques should tremendously improve the precision of elicited preferences (see Soumalias et al., 2024).
After taking in students’ reported preferences, the algorithm used to find an equilibrium needs to be able to solve maximization problems that find each student’s most-preferred course schedule subject to her budget and priority-specific prices. With thousands of courses, overlapping lecture times and prerequisite requirements create considerable scheduling constraints that students should not be expected to handle through reported preferences alone. To deal with these constraints, the maximization problem should prevent students from obtaining seats in courses that meet at the same time. Relatedly, it should accommodate course prerequisite constraints, blocking students from enrolling in courses they cannot obtain. As many degree programs require students to take certain courses in a given term, making these assignments outside the mechanism would reduce preference reporting mistakes. These modifications would make our mechanism simpler to implement in practice.
@@
Finding a Pseudo-Market Equilibrium with Priorities with students’ reported preferences is a computationally intense problem. Budish et al. (2017) estimate that 20 times the computational time used for business school students at Wharton would be necessary to implement the A-CEEI mechanism for the full student body at Ohio State University. However, equilibrium computation for the PMP mechanism may be less challenging than this estimate suggests. Students in unrelated programs often have limited overlap in course interests. The structure of priority-specific prices simplifies the problem of finding a student’s demand, as the price is zero or unaffordable at all but one of a course’s priority levels. In our simulations, the average timing of the PMP mechanism improves on that of the A-CEEI mechanism by a factor of two (see Appendix D.2). Still, incorporating scheduling constraints and preference complementarities may be challenging and should be carefully addressed in practice.
@@ Appendix A Appendix: Proofs
$p_{c,r}(t)\in\begin{cases}\left\{0\right\}&r>r^{*}_{c}(t)\\ {}[0,\overline{b})&r=r^{*}_{c}(t)\\ [\overline{b},+\infty)&r<r^{*}_{c}(t)\end{cases}.$ (A2)
We will consider an auxiliary enlargement of this set, $\tilde{{\cal T}}=[-1,R\overline{b}+1]^{M}$ and define $p_{c,r}(\tilde{t})$ using equation (A1) for $\tilde{t}\in\tilde{\cal T}$. For each $s\in{\cal S}$, define the demand function $d_{s}:\tilde{{\cal T}}\rightarrow 2^{\cal C}$ by
@@
where the $\tau_{s,x_{s}}$ are student- and schedule-specific taxes chosen to ensure that the demand is single-valued. For each course $c\in{\cal C}$, the excess demand function $z_{c}:\tilde{{\cal T}}\rightarrow\mathbb{Z}$ is defined by
@@
where $x^{*}_{s}=d_{s}(t^{*})$ for all $s\in{\cal S}$. Excess demand is bounded, as $-N\leq z_{c}\leq N-1$ for all $c\in{\cal C}$. We also define a budget surface for each student $s\in{\cal S}$ and schedule $x_{s}\subseteq{\cal C}$ as
@@
Next, we define a truncation function trunc$:\tilde{{\cal T}}\rightarrow{\cal T}$, where for each $c\in{\cal C}$,
@@
$f(\tilde{t})=\text{trunc}(\tilde{t})+\gamma z(\text{trunc}(\tilde{t})).$
Suppose that $f$ has a fixed point $\tilde{t}^{*}=f(\tilde{t}^{*}),$ and denote its truncation by $t^{*}=\text{trunc}(\tilde{t}^{*})$. We show that the prices $\{p_{c,r}(t^{*})\}_{c\in{\cal C},{r\in\{1,...,R\}}}$ defined by equation (A1), allocation $x^{*}_{s}=d_{s}(t^{*})$, and budgets $b^{*}_{s}=b_{s}+\tau_{s,x^{*}_{s}}$ for all $s\in{\cal S}$ constitute an exact competitive equilibrium (or ($0,\beta$)-Pseudo-Market Equilibrium with Priorities as in Definition 4). We might only need to slightly adjust $t^{*}$ to find another fixed point that satisfies condition $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(t^{*})\}}x^{*}_{s,c}>0$.
@@
  • To make sure the condition $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(t^{*})\}}x^{*}_{s,c}>0$ is satisfied for any course $c$ with $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ (see condition ($\ast$2)), assume from the contrary that course $c$ has demand $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ but demand from students at the cutoff level of priority is 0: $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(t^{*})\}}x^{*}_{s,c}=0$. Then, we can consider $\hat{t}_{c}^{*}=t^{*}_{c}+\overline{b}-p_{c,r_{c}^{*}(t^{*})}$, where the students at the cutoff level of priority a face of price $\overline{b}$ and the prices at lower priority levels do not change. With $\hat{t}^{*}=(\hat{t}^{*}_{c},t^{*}_{-c})$, the cutoff increases to level $r^{*}_{c}(\hat{t}^{*})=r^{*}_{c}(t^{*})+1$, the prices become $p_{c,r^{*}_{c}(\hat{t}^{*})}=0$ and ${p}_{c,r^{*}_{c}(t^{*})}=\overline{b}$, and the adjustment does not change the demand for any student or excess demand for any course. If we still have $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(\hat{t}^{*})\}}x^{*}_{s,c}=0$, we repeat the price adjustment until the condition is satisfied.
    Finally, $z_{c}(\hat{t}^{*})=z_{c}(t^{*})=0$ implies $\hat{t}^{*}_{c}<R\overline{b}$ and $\hat{t}^{*}_{c}>t^{*}_{c}\geq 0$. Therefore, $\hat{t}^{*}_{c}=\text{trunc}(\hat{t}^{*}_{c})+z_{c}(\hat{t}^{*})$, meaning $\hat{t}^{*}$ satisfies both the fixed point equation and $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(\hat{t}^{*})\}}x^{*}_{s,c}>0$.
Overall, if $f$ has a fixed point $\tilde{t}^{*}=f(\tilde{t}^{*})$, its truncation $t^{*}=\text{trunc}(\tilde{t}^{*})$ or adjustment $\hat{t}^{*}$ is a competitive equilibrium price vector for allocation $x^{*}_{s}=d_{s}(t^{*})$ and budgets $b_{s}^{*}$ for all $s\in{\cal S}$. Though, the fixed point of operator $f$ might fail to exist. Following Budish (2011), we define a “convexification” of $f$, $F:\tilde{\cal T}\rightarrow\tilde{\cal T}$, by
$F(\tilde{t})=co\{y:\exists\text{ a sequence }\tilde{t}^{w}\rightarrow\tilde{t},\tilde{t}\neq\tilde{t}^{w}\in\tilde{\cal T}\text{ such that }f(\tilde{t}^{w})\rightarrow y\},$
where $co$ denotes the convex hull of the set. $F$ is nonempty and convex and $\tilde{\cal T}$ is compact and convex. From Lemma 2.4 of Cromme and Diener (1991), $F$ is an upper hemicontinuous correspondence and hence has a fixed point by Kakutani’s fixed point theorem. We denote a fixed point by $\tilde{t}^{*}\in F(\tilde{t}^{*})$, and let again $t^{*}=\text{trunc}(\tilde{t}^{*})$ be its truncation.
The remainder of the proof uses $F$ to show that there exists a vector $t^{*}$ with excess demand at most a distance of $\sqrt{kM/2}$ from a perfect market-clearing excess demand vector. The reduction of the $M\!R$-dimensional space of prices $\{p_{c,r}\}_{c\in{\cal C},{r\in{\{1,...,R\}}}}$ to $M$-dimensional space $\tilde{{\cal T}}=[-1,R\overline{b}+1]^{M}$ allows us to use the steps of Budish (2011) to establish the existence of a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities. For completeness, we provide adapted steps in Appendix B.∎
@@
The final term must equal zero, as allocations $y$ and $x^{*}$ assign the same number of seats in each course, Also, $p^{*}_{c,r}-p^{*}_{c,r-1}\leq 0$ for all levels $r$, and we must have $\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})<0$ for some course $c$ and level $r$. This implies $y_{c}\nsucceq_{c}x^{*}_{c}$ and confirms that allocation $x^{*}$ is approximately priority-constrained efficient.∎
@@
where $\Phi_{s}^{N}(v_{s},v_{-s})$ denotes the course schedule obtained by student $s$ when she reports type $v_{s}$ and all other students report $v_{-s}$. $Pr(v_{-s}|v_{-s}\sim iid(m))$ denotes the probability that profile $v_{-s}$ is realized when other students’ types $v_{-s}$ are independent and identically distributed according to $m\in\Delta{\cal V}$. In other words, $\phi^{N}(v_{s},m)$ describes the random outcome that student $s$ expects to receive when she reports $v_{s}$ and the other students’ types are distributed independently and identically according to $m$—and they report their types truthfully. Using this notation, we have the following definition.
@@
[…]
@@ Appendix B Proof of Theorem 1 (Omitted Details)
    Taxes are small: $-\varepsilon<\tau_{s,x_{s}}<\varepsilon$.
@@
Fix an allocation $x=(x_{s})_{s\in{\cal S}}$. Budish (2011) shows that it is possible to choose $\left\{\tau_{s,x_{s}}\right\}_{s\in{\cal S},x_{s}\subseteq{\cal C}}$ that satisfy the first four conditions. We now establish that it is always possible to slightly change taxes such that condition (v) is also satisfied. Let us assume that more than $M$ budget constraints $H(s,x_{s})$ intersect and denote
=${\cal I}=\left\{s\in{\cal S}:\cap_{s}H(s,x_{s})\neq\varnothing\right\},$
with $|{\cal I}|>M.$ For each $\tilde{t}\in\tilde{\cal T}$, consider prices $\left\{p_{c,r}(\tilde{t})\right\}_{c\in{\cal C},r\in{\{1,...,R\}}}$ defined by equation (A1) and the cutoffs defined by (A2). The definition of cutoffs $r^{*}_{c}(\tilde{t})$ implies that
@@
In addition, the definition of cutoffs implies $p_{c,r_{s,c}}\geq\overline{b}$ for $1\leq r_{s,c}<r^{*}_{c}(\tilde{t})$. Therefore, a seat in course $c$ is not allocated to agent $s$ for $r_{s,c}<r^{*}_{c}(\tilde{t})$; that is, $x_{s,c}=0.$ Hence, we also have
@@
Therefore, we obtain that $x_{s,c}\cdot p_{c,r_{s,c}}(\tilde{t})$ might be non-zero only if $r_{s,c}=r^{*}_{c}(\tilde{t})$. That is, entries in agent $s$’s budget constraint, $s\in{\cal I}$, are non-zero only if $r_{s,c}=r^{*}_{c}(\tilde{t})$. Denote $p_{c}(\tilde{t})\equiv p_{r^{*}_{c}}(\tilde{t})$. Hence, we obtain that the set of equations $\left\{s\in{\cal I}:H(s,x_{s})=0\right\}$ is the set of linear equations with coefficients $x_{s,c}\in\{0,1\}$:
$\left\{\begin{array}[]{cc}x_{s,1}\cdot p_{1}(\tilde{t})+x_{s,2}\cdot p_{2}(\tilde{t})+...+x_{s,M}\cdot p_{M}(\tilde{t})&=b_{s}+\tau_{s,x_{s}}\\ ....&....\\ x_{s^{\prime},1}\cdot p_{1}(\tilde{t})+x_{s^{\prime},2}\cdot p_{2}(\tilde{t})+...+x_{s^{\prime},M}\cdot p_{M}(\tilde{t})&=b_{s^{\prime}}+\tau_{s^{\prime}x}\end{array}\right.$
for $s,s^{\prime}\in{\cal I}$. Since $|{\cal I}|>M$ for any $\tilde{t}\in{\cal T}$, there are at most $M$ independent linear equations of prices. So, the Rouché–Capelli theorem implies that we can choose $\tau_{s,x_{s}}$ such that only at most $M$ equations are satisfied for any $\tilde{t}$. [^29] Note that when we change $\tau_{s,x_{s}}$, some other budget constraints might start intersecting. Since the set of possible intersecting budget constraints is finite and $\tau_{s,x_{s}}$ varies continuously, we can always choose $\tau_{s,x_{s}}$ without influencing the intersection property of the other budget constraints.
@@
  • If $t^{*}=\tilde{t}^{*}$, then $F(\tilde{t}^{*})=F(t^{*})=f(t^{*})$ and, thus, $t^{*}=\tilde{t}^{*}\in F(\tilde{t}^{*})=f(\tilde{t}^{*})$. Therefore, $t^{*}$ is a fixed point. Hence, as we established earlier, it is an exact competitive equilibrium price vector.
@@
    For any $\tilde{t}\in\tilde{\cal T}\setminus{\cal T}$, (i) $f(\tilde{t})=f($$(\tilde{t}))$ and (ii) $F(\tilde{t})\subseteq F($$(\tilde{t}))$.
@@
    As a result, since $F(t^{*})=f(t^{*})$ and $\tilde{t}^{*}\in F(\tilde{t}^{*})$, $\tilde{t}^{*}\in F(t^{*})=f(t^{*})=f(\tilde{t}^{*})$ and, thus, $\tilde{t}^{*}=f(\tilde{t}^{*})$. Therefore, $t^{*}$ is an exact competitive equilibrium price vector.
@@
    Each $t^{\phi}$ is on the “affordable” side of the $\ell$th budget constraint if $\phi_{\ell}=0$ and is on the “unaffordable” side if $\phi_{\ell}=1$.
@@
$H^{0}_{\ell}$ =$=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)\leq b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
$H^{1}_{\ell}$ =$=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)>b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
@@
Step 6. We now show that a perfect market-clearing excess demand vector lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$. For this purpose, we first show that for any $y\in F(t^{*})$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ Take some $y\in F(t^{*})$. Consider a sequence $t^{w}\rightarrow t^{*},t^{*}\neq t^{w}\in\tilde{\cal T}\text{ such that }f(t^{w})\rightarrow y^{\prime}$. Note that the sequence $t^{w}$ consists of a finite number of subsequences $t^{w,\phi}\in\pi(\phi)\cap B_{\delta}(t^{*})$ for some $\phi\in\Phi$ [^30] Some subsequences can have only a finite number of elements. Since all elements of the subsequence $t^{w,\phi}$ are on the same side of set $H^{\phi}_{\ell}$, every agent has the same choice at every point. Hence, if the subsequence has an infinite number of elements, we have $t^{w,\phi}\rightarrow t^{*}$ and
@@
Therefore, the limit of $f(t^{w})$ for the original sequence $t^{w}$ must be also $t^{*}+\gamma z(t^{\phi^{\prime}})$ for some $\phi^{\prime}\in\Phi$. So, $y^{\prime}=t^{*}+\gamma z(t^{\phi^{\prime}})$. Since, by definition, $y$ is a convex combination of such $y^{\prime}$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$
@@
$\tilde{t}^{*}=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi}).$
for some $\{\lambda_{\phi}\}_{\phi\in\Phi}$ with $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ We also denote
@@
Steps 7-9. The structure of excess demand has the same geometric structure as in Budish (2011). In particular, let $L^{\prime}$ be the number of agents whose budget constraints intersect at price $t^{*}$, and renumber agents so that $s=1,...,L^{\prime}$. We denote the number of budgets of student $s$ that intersect at $t^{*}$ as $w_{s}$. Since at most $M$ budgets constraints can intersect, we must have $L\equiv\sum_{s=1}^{L^{\prime}}w_{s}\leq M$. We also denote the bundles pertaining to $s$’s budget constraints as $x_{s}^{1}\succ...\succ x_{s}^{w_{s}}$.

### v2

@@ 4 Discussion
This paper explores undergraduate course allocation as a many-to-many matching problem. We design a deterministic mechanism, the Pseudo-Market with Priorities (PMP) mechanism, which extends the approximate competitive equilibrium from equal incomes mechanism of Budish (2011) to accommodate course priorities and has the same worst-case bound on the market-clearing error. This mechanism satisfies approximate notions of stability, efficiency, fairness, and strategy-proofness. Our empirical findings exhibit that it outperforms an idealized version of the mechanism used in practice in student satisfaction and fairness.
The gains from introducing the PMP mechanism will generally depend on university-specific factors. When a large set of equally-prioritized students compete for a small set of seats across courses, the commonly-used RSD mechanism produces allocations with significant envy. Instead, the PMP mechanism bounds envy among these students. Hence, the fairness improvements from introducing our mechanism will be large in environments with many over-demanded courses. Similarly, if a university uses a coarser priority structure, there will be more equally-prioritized students, increasing the gains from introducing the PMP mechanism. In contrast, for universities that employ finer priority structures (e.g., prioritizing students based on units completed or grade point average), the PMP and RSD mechanisms will result in similar outcomes.
@@
A university’s practical implementation of the PMP mechanism will also play a role in its benefits. Our model assumes students can report preferences over all possible course schedules. In practice, reporting preferences requires a language simple enough for students to use and expressive enough to elicit students’ actual preferences. The implementation of the A-CEEI mechanism in Budish et al. (2017) allows students to report cardinal values for individual courses and adjustments for pairs of courses. Changes to this language are necessary to make it accessible to undergraduate students. In particular, it should accommodate overlapping lecture times, prerequisite constraints, required courses, and credit limits, blocking students from receiving unacceptable schedules. The deployment of machine learning-based techniques may also improve the precision of elicited preferences (Soumalias et al., 2024).
@@
Finding an equilibrium given reported preferences is computationally intense. Budish et al. (2017) estimate that 20 times the time used at Wharton would be necessary to implement the A-CEEI mechanism for the full student body at Ohio State University. However, implementing the PMP mechanism may be less challenging than this estimate suggests. Since a course’s price is zero or unaffordable at all but one of its priority levels, students face few courses with a positive but affordable price. In our simulations, with additive student utilities, the average timing of the PMP mechanism improves on the A-CEEI mechanism by a factor of two (see Appendix D.2). Still, incorporating scheduling constraints and preference complementarities may be challenging and should be carefully addressed in practice.
Finally, we note that our analysis applies to many other matching problems. Two under-explored examples are assigning teaching assistants to courses and referees to soccer matches. Extending our results to these and other applications is left for future research.
@@ Appendix A Appendix: Proofs
$p_{c,r}(t)\in\begin{cases}\left\{0\right\}&r>r^{*}_{c}(t)\\ {}[0,\overline{b})&r=r^{*}_{c}(t)\\ [\overline{b},R\overline{b}]&r<r^{*}_{c}(t)\end{cases}.$ (A2)
[…]
@@
where the $\tau_{s,x_{s}}$ are student- and schedule-specific taxes chosen to ensure that the demand is single-valued. Taxes are chosen to favor more-preferred bundles (i.e., if $x^{\prime}_{s}\succ_{s}x_{s}$, then $\tau_{s,x^{\prime}_{s}}>\tau_{s,x_{s}}$). For each course $c\in{\cal C}$, the excess demand function $z_{c}:\tilde{{\cal T}}\rightarrow\mathbb{Z}$ is defined by
@@
where $x^{*}_{s}=d_{s}(t^{*})$ for all $s\in{\cal S}$. Excess demand is bounded, as $N\leq z_{c}\leq N-1$ for all $c\in{\cal C}$. We also define a budget surface for each student $s\in{\cal S}$ and schedule $x_{s}\subseteq{\cal C}$ as
@@
[…]
@@
$).f(\tilde{t})=\text{trunc}(\tilde{t})+\gamma z(\text{trunc}(\tilde{t}$)).
[…]
@@
[…]
The remainder of the proof uses $F$ to show that there exists budgets for which excess demand at prices $t^{*}$ are at most a distance of $\sqrt{kM/2}$ from a perfect market-clearing excess demand vector. The reduction of the MRM\!R-dimensional space of prices $\{p_{c,r}\}_{c\in{\cal C},{r\in{\{1,...,R\}}}}$ to $M$-dimensional space $\tilde{{\cal T}}=[-1,R\overline{b}+1]^{M}$ allows us to use the steps of Budish (2011) to establish the existence of a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities. For completeness, we provide adapted steps in Appendix B.∎
@@
The final term must equal zero, as allocations $y$ and $x^{*}$ assign the same number of seats in each course. Also, $p^{*}_{c,r}-p^{*}_{c,r-1}\leq 0$ for all levels $r$, so we must have $\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})<0$ for some course $c$ and level $r$. This implies $y_{c}\nsucceq_{c}x^{*}_{c}$ and confirms that allocation $x^{*}$ is approximately priority-constrained efficient.∎
@@
where $\Phi_{s}^{N}(v_{s},v_{-s})$ denotes the course schedule obtained by student $s$ when she reports type $v_{s}$ and all other students report $v_{-s}$. $Pr(v_{-s}|v_{-s}\sim iid(m))$ denotes the probability that profile $v_{-s}$ is realized when other students’ types $v_{-s}$ are independent and identically distributed according to $m\in\Delta{\cal V}$. In other words, $\phi^{N}(v_{s},m)$ describes the random outcome that student $s$ expects to receive when she reports $v_{s}$ and the other students’ types are distributed independently and identically according to mm—and they report their types truthfully. Using this notation, we have the following definition.
@@
A semi-anonymous direct mechanism $\{\Phi^{N}\}_{N\in{\cal N}}$ is envy-free but for tie-breaking if for each $N$ there exists $x^{N}:({\cal V}\times[0,1])^{N}\rightarrow\Delta(X_{0}^{N})$, symmetric over its coordinates, such that $\Phi^{N}(v)=\int_{\ell\in[0,1]^{N}}x^{N}(v,\ell)d\ell,$ and, for all $s,s^{\prime},N,v$, and $\ell$, if $\ell_{s}\geq\ell_{s^{\prime}}$, and $v_{s}$ and $v_{s^{\prime}}$ belong to the same group $S_{g}$ (that is, $r_{s}=r_{s^{\prime}}$), then $u_{v_{s}}(x^{N}_{s}(v,\ell))\geq u_{v_{s}}(x^{N}_{s^{\prime}}(v,\ell))$.
The PMP mechanism assigns each student a budget using a uniformly distributed lottery $\ell_{s}\sim U[0,1]$ as $b_{s}=1+\ell_{s}\beta$. Because budgets are random, the PMP assignment for each $v\in{\cal V}^{N}$ can be represented as $\Phi^{N}(v)=\int_{\ell\in[0,1]^{N}}x^{N}(v,\ell)d\ell,$ where $x^{N}(v,\ell)\in\{0,1\}^{MN}$ assigns a course schedule to each student in ${\cal S}^{N}.$ Definition 4 guarantees that student $s$ with budget $b_{s}$ (lottery number $l_{s}$) prefers her course schedule to the course schedule of student $s^{\prime}$ in the same group with budget $b_{s^{\prime}}<b_{s}$ ($\ell_{s^{\prime}}<\ell_{s}$). This confirms that the PMP mechanism is strategy-proof in the large and completes the proof.∎
@@ Appendix B Proof of Theorem 1 (Omitted Details)
    Taxes are small: $\varepsilon<\tau_{s,x_{s}}<\varepsilon$.
@@
$\forall s\in{\cal I},c\in{\cal C}:r_{s,c}>r^{*}_{c}(\tilde{t}),\hskip 8.5359ptx_{s,c}\cdot p_{c,r_{s,c}}(\tilde{t})=0.$
[…]
$\forall s\in{\cal I},c\in{\cal C}:r_{s,c}<r^{*}_{c}(\tilde{t}),\hskip 8.5359ptx_{s,c}\cdot p_{c,r_{s,c}}(\tilde{t})=0.$
[…]
$\left\{\begin{array}[]{cc}x_{s,1}\cdot p_{1,r_{s,1}}(\tilde{t})+x_{s,2}\cdot p_{2,r_{s,2}}(\tilde{t})+...+x_{s,M}\cdot p_{M,r_{s,M}}(\tilde{t})&=b_{s}+\tau_{s,x_{s}}\\ ....&....\\ x_{s^{\prime},1}\cdot p_{1,r_{s^{\prime},1}}(\tilde{t})+x_{s^{\prime},2}\cdot p_{2,r_{s^{\prime},2}}(\tilde{t})+...+x_{s^{\prime},M}\cdot p_{M,r_{s^{\prime},M}}(\tilde{t})&=b_{s^{\prime}}+\tau_{s,x^{\prime}_{s}}\end{array}\right$.
Here, $s,s^{\prime}\in{\cal I}$, the coefficients $x_{s,c}\in\{0,1\}$ are in line with the bundles $x_{s}$ on the budget constraints, and each price satisfies $p_{c,r_{s,c}}=p_{c,r^{*}_{c}}$ or is such that $x_{s,c}\cdot p_{c,r_{s,c}}=0$. Equivalently, setting $\tilde{x}_{s,c}=1$ if $x_{s,c}=1$ and $r_{s,c}=r^{*}_{c}$ and $\tilde{x}_{s,c}=0$, we have the following set of equations:
$\left\{\begin{array}[]{cc}\tilde{x}_{s,1}\cdot p_{1,r^{*}_{1}}(\tilde{t})+\tilde{x}_{s,2}\cdot p_{2,r^{*}_{2}}(\tilde{t})+...+\tilde{x}_{s,M}\cdot p_{M,r^{*}_{M}}(\tilde{t})&=b_{s}+\tau_{s,x_{s}}\\ ....&....\\ \tilde{x}_{s^{\prime},1}\cdot p_{1,r^{*}_{1}}(\tilde{t})+\tilde{x}_{s^{\prime},2}\cdot p_{2,r^{*}_{2}}(\tilde{t})+...+\tilde{x}_{s^{\prime},M}\cdot p_{M,r^{*}_{M}}(\tilde{t})&=b_{s^{\prime}}+\tau_{s,x^{\prime}_{s}}\end{array}\right$.
[…]
@@
[…]
@@
    For any $\tilde{t}\in\tilde{\cal T}\setminus{\cal T}$, (i) $f(\tilde{t})=f($trunc(t~))($\tilde${t})) and (ii) $F(\tilde{t})\subseteq F($trunc(t~))($\tilde${t})).
@@
[…]
@@
[…]
@@
$H^{0}_{\ell}$ $=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)\leq b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
$H^{1}_{\ell}$ $=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)>b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
@@
Step 6. We now show that a perfect market-clearing excess demand vector lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$. For this purpose, we first show that for any $y\in F(t^{*})$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}\gamma z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ Take some $y\in F(t^{*})$. Consider a sequence $t^{w}\rightarrow t^{*},t^{*}\neq t^{w}\in\tilde{\cal T}\text{ such that }f(t^{w})\rightarrow y^{\prime}$. Note that the sequence $t^{w}$ consists of a finite number of subsequences $t^{w,\phi}\in\pi(\phi)\cap B_{\delta}(t^{*})$ for some $\phi\in\Phi$. [^27] Some subsequences can have only a finite number of elements. Since all elements of the subsequence $t^{w,\phi}$ are on the same side of set $H^{\phi}_{\ell}$, every agent has the same choice at every point. Hence, if the subsequence has an infinite number of elements, we have $t^{w,\phi}\rightarrow t^{*}$ and
@@
Therefore, the limit of $f(t^{w})$ for the original sequence $t^{w}$ must be also $t^{*}+\gamma z(t^{\phi^{\prime}})$ for some $\phi^{\prime}\in\Phi$. So, $y^{\prime}=t^{*}+\gamma z(t^{\phi^{\prime}})$. Since, by definition, $y$ is a convex combination of such $y^{\prime}$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}\gamma z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$
@@
$\tilde{t}^{*}=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}\gamma z(t^{\phi}).$
for nonnegative $\{\lambda_{\phi}\}_{\phi\in\Phi}$ with $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ We also denote
@@
Steps 7-8. The structure of excess demand has the same geometric structure as in Budish (2011). In particular, let $L^{\prime}$ be the number of agents whose budget constraints intersect at price $t^{*}$, and renumber agents so that $s=1,...,L^{\prime}$. We denote the number of budgets of student $s$ that intersect at $t^{*}$ as $w_{s}$. Since at most $M$ budgets constraints can intersect, we must have $L\equiv\sum_{s=1}^{L^{\prime}}w_{s}\leq M$. We also denote the bundles pertaining to $s$’s budget constraints as $x_{s}^{1}\succ...\succ x_{s}^{w_{s}}$.